### v3

@@ Abstract
Techniques to autonomously drive research have been prominent in Computational Scientific Discovery, while Synthetic Biology is a field of science that focuses on designing and constructing new biological systems for useful purposes. Here we seek to apply logic-based machine learning techniques to facilitate cellular engineering and drive biological discovery. Comprehensive databases of metabolic processes called genome-scale metabolic network models (GEMs) are often used to evaluate cellular engineering strategies to optimise target compound production. However, predicted host behaviours are not always correctly described by GEMs, often due to errors in the models. The task of learning the intricate genetic interactions within GEMs presents computational and empirical challenges. To address these, we describe a novel approach called Boolean Matrix Logic Programming (BMLP) by leveraging boolean matrices to evaluate large logic programs. We introduce a new system, $BMLP_{active}$, which efficiently explores the genomic hypothesis space by guiding informative experimentation through active learning. In contrast to sub-symbolic methods, $BMLP_{active}$ encodes a state-of-the-art GEM of a widely accepted bacterial host in an interpretable and logical representation using datalog logic programs. Notably, $BMLP_{active}$ can successfully learn the interaction between a gene pair with fewer training examples than random experimentation, overcoming the increase in experimental design space. $BMLP_{active}$ enables rapid optimisation of metabolic models to reliably engineer biological systems for producing useful compounds. It offers a realistic approach to creating a self-driving lab for microbial engineering.
@@ 1 Introduction
A priority in scientific research is the development of autonomous systems to enhance the scientific discovery process. Currently, most efforts in this domain focus on either laboratory automation or machine learning applied to scientific datasets. Sub-symbolic approaches usually demand a large number of annotated data and extensive trial-and-error. It is often difficult to modify the structure of these machine learning systems or extract human-interpretable information. The application of these systems in an experimental workflow is limited by data efficiency, flexibility and interpretability. There is a resurgence of interest in incorporating AI in experimental automation and scientific understanding simultaneously. In this work, we investigate an integration of techniques for planning experiments, optimising resource utilisation and interpretable machine learning to support biological discovery.
@@
In the pursuit of understanding natural phenomena, scientists often formulate plausible theories and design experiments to test competing hypotheses. This process typically involves a careful selection of experimental conditions and the logical inference mechanism of abduction. Experimentation can be optimised through active learning, where computational methods are employed to strategically select informative training examples to combat resource limitations. On the other hand, Inductive Logic Programming (ILP) ILP1991 offers an automated approach to abduction, representing observations, hypotheses and background knowledge through interpretable logic programs. In ILP, hypotheses are learned to explain observational data with respect to the background knowledge. “Askable” hypotheses, known as abducibles, that might explain contradictions between observations and known background knowledge are first assumed to hold ase_progol . The resulting extended knowledge, enriched with abduced hypotheses, is subsequently verified against observations TCIE .
@@
Our study explores the applicability of abductive reasoning and active learning to comprehensive biological systems. Given that biological relationships are commonly described logically, ILP is particularly adept at operating on biological knowledge bases. The integration of abductive reasoning and active learning via ILP was successfully demonstrated in the context of biological discovery by the prominent Robot Scientist King04:RobotScientist . The Robot Scientist performed active learning by strategically selecting key experiments to achieve more data- and cost-effective gene function learning than random experimentation. However, this demonstration was limited to only 17 genes in the aromatic amino acid pathway of yeast. In the present study, we advance the scope of logic-based gene function learning by looking at Escherichia coli (E. coli). The genome-scale metabolic model (GEM), iML1515 iML1515 , encompasses 1515 genes and 2719 metabolic reactions of the E. coli strain K-12 MG1655, a versatile host organism for metabolic engineering to produce specific compounds. Notably, the iML1515 model represents a 90-fold increase in complexity compared to the aromatic amino acid pathway investigated by the Robot Scientist.
Despite the extensive mapping of the E. coli genome to metabolic functions in iML1515, it was shown to yield inaccurate phenotype predictions bernstein_evaluating_2023 . A classical method to address errors in metabolic pathways involves auxotrophic growth experiments, where specific genes are deleted to render the organism incapable of synthesizing essential compounds beadle_genetic_1941 . For accurate simulation of auxotrophic mutant experiments, the computational representation of GEMs must faithfully depict the underlying network structure and effectively facilitate the exploration of metabolic pathways.
Considering these requirements, Petri nets emerge as suitable modelling tools for GEMs. Petri nets are a class of directed bipartite graphs with diverse applications in Computer Science and Biology reisig_understanding_2013 ; sahu_advances_2021 . Petri nets can represent a structured knowledge base while capturing dynamic behaviours in a system. In contrast, logic programs allow reasoning over relationships between entities but cannot directly reflect the changes in concepts and entity states. To integrate logic reasoning with Petri nets, we explore a more holistic approach to learning metabolic networks that combines recursive logic programs with pathway reachability analysis.
@@
Figure 1: The hypotheses and instances are represented as logic programs to describe gene functions (gene-enzyme associations) and auxotrophic mutant experiment conditions respectively. $BMLP_{active}$ encodes the GEM iML1515 and simulates the auxotrophic mutant experiment outcomes. It actively consults a data source to request ground truth labels that minimise the expected experimental cost based on a user-provided cost function when pruning hypotheses. The underlying BMLP iteratively refutes gene function hypotheses inconsistent with labelled training examples. While $BMLP_{active}$ to date only learns from synthetic data, a laboratory robot can be integrated to perform selected experiments.
@@
Learning complex biological systems like GEMs presents significant challenges in terms of computational complexity and empirical feasibility. To address these challenges, we present a new system $BMLP_{active}$ (Figure 1) based on a novel approach called Boolean Matrix Logic Programming (BMLP) and an implementation in SWI-Prolog. We use logic programs to encode the biochemical, genetic, and genomic relationships in iML1515. While simulating auxotrophic mutant experiments, logical inferences to predict phenotypic effects are accelerated by boolean matrices via BMLP. This speeds up the runtime for predicting phenotypic effects by a factor of 170 compared to SWI-Prolog without BMLP on a single computation thread. We implement active learning to select auxotrophic mutant experiments that minimise the expected experimental cost via a user-defined cost function. Empirically, we delete gene functions in iML1515 and use $BMLP_{active}$ to learn hypotheses to recover the deleted gene functions. $BMLP_{active}$ successfully restores these removed gene functions through active learning while reducing 90% the optional nutrient substance cost required by randomly selected experiments. It additionally lowers the number of experimental data needed to recover these gene functions compared to random experiment selection, reducing the sample complexity. When operating within a finite budget, our approach can deliver optimal experimental outcomes, whereas studies employing random experiment selection might not reach completion.
@@
Furthermore, we apply $BMLP_{active}$ to learning digenic functions, which holds significant implications for drug development and therapeutic interventions costanzo_global_2016 . These digenic interactions are related to isoenzymes - two enzymes responsible for the same reaction. The intricate relationship between complex genotypes and phenotypes remains largely unexplored in most organisms costanzo_global_2019 . Additionally, some digenic interactions are dynamic and can vary depending on the growth conditions bernstein_evaluating_2023 . A comprehensive understanding of digenic interactions would expedite strain engineering efforts. We show that $BMLP_{active}$ can converge to the correct gene-isoenzyme mapping with as little as 20 training examples. This shows a significant promise for $BMLP_{active}$ to address complex genetic interactions on the whole genome level.
Our contributions are:
  1. 1.
    We describe a novel problem of Boolean Matrix Logic Programming (BMLP).
  2. 2.
    We implement a BMLP algorithm, BMLP-IE, in SWI-Prolog to predict phenotypic effects by computing reachability queries in GEMs based on a subclass of Petri net, which empirically improves simulation runtime by 170-fold compared to SWI-Prolog without BMLP-IE.
  3. 3.
    We built a BMLP system, $BMLP_{active}$, which uses BMLP-IE to actively select experiment instances for learning gene functions in a GEM.
  4. 4.
    We present a theoretical sample complexity ratio between active learning and random experiment sampling and empirically highlight the effect of actively selecting examples in reducing the total training examples needed.
  5. 5.
    We empirically show that $BMLP_{active}$ reduces 90% optional nutrient substance cost and the number of experiments needed for learning gene functions compared to randomly sampled experiments.
  6. 6.
    Our empirical results demonstrate that $BMLP_{active}$ can be devised to learn gene-isoenzyme associations in GEM with a faster convergence to the correct gene-isoenzyme mapping than randomly sampled experiments.
The rest of this paper is organised as follows. In Section 2, we summarise work on machine learning of genome-scale metabolic networks, matrix algebra, Petri nets and active learning. In Section 3, we introduce key concepts in matrix algebra and an elementary class of Petri nets that we use to model a GEM, called one-bounded elementary nets (OENs). In Section 4, we describe a transformation from an OEN to a subclass of recursive logic programs, linear and immediately recursive datalog. In Section 5, we introduce our BMLP framework to evaluate this class of logic programs. In Section 6, we describe our active learning approach based on prior work ase_progol ; progol . In Section 7, we explain the $BMLP_{active}$ implementation for encoding iML1515, GEM simulation and active learning gene functions. Section 8 includes our empirical results on simulation runtime, phenotype prediction, active learning gene functions and gene-isoenzyme mapping. In Section 9, we explain the current limitations of $BMLP_{active}$ and discuss future work.

### v4

@@ Abstract
Reasoning about hypotheses and updating knowledge through empirical observations are central to scientific discovery. In this work, we applied logic-based machine learning methods to drive biological discovery by guiding experimentation. Genome-scale metabolic network models (GEMs) - comprehensive representations of metabolic genes and reactions - are widely used to evaluate genetic engineering of biological systems. However, GEMs often fail to accurately predict the behaviour of genetically engineered cells, primarily due to incomplete annotations of gene interactions. The task of learning the intricate genetic interactions within GEMs presents computational and empirical challenges. To efficiently predict using GEM, we describe a novel approach called Boolean Matrix Logic Programming (BMLP) by leveraging Boolean matrices to evaluate large logic programs. We developed a new system, $BMLP_{active}$, which guides cost-effective experimentation and uses interpretable logic programs to encode a state-of-the-art GEM of a model bacterial organism. Notably, $BMLP_{active}$ successfully learned the interaction between a gene pair with fewer training examples than random experimentation, overcoming the increase in experimental design space. $BMLP_{active}$ enables rapid optimisation of metabolic models to reliably engineer biological systems for producing useful compounds. It offers a realistic approach to creating a self-driving lab for biological discovery, which would then facilitate microbial engineering for practical applications.
@@ 1 Introduction
In the pursuit of understanding natural phenomena, scientists often formulate plausible theories and design experiments to test competing hypotheses. This process typically involves careful experimental design and abductive reasoning. Experimentation can be optimised through active learning mitchell_generalization_1982 ; cohn_improving_1994 , where informative training examples are strategically selected to combat resource limitations. Inductive Logic Programming (ILP) ILP1991 offers an automated approach to abduction, representing observations, hypotheses and background knowledge through interpretable logic programs. In ILP, hypotheses are learned to explain observational data with respect to the background knowledge. “Askable” hypotheses, known as abducibles, that might explain contradictions between observations and prior background knowledge are proposed ase_progol ; TCIE . The extended knowledge, enriched with abduced hypotheses, is subsequently verified against observations.
@@
Our work explored the applicability of abductive reasoning and active learning to identify the functions of metabolic genes in model bacterial organisms. Given that biological relationships are commonly described logically, ILP is particularly adept at operating on biological knowledge bases. The integration of abductive reasoning and active learning via ILP was successfully demonstrated in the context of biological discovery by the Robot Scientist King04:RobotScientist . The Robot Scientist performed active learning by strategically selecting key experiments to achieve more data- and cost-effective gene function learning than random experimentation. However, this demonstration was limited to only 17 genes in the aromatic amino acid pathway of yeast. Our work significantly advanced the application of this integrative learning paradigm to genome-scale biological discoveries by examining genome-scale metabolic models (GEMs). We examined the GEM model iML1515 iML1515 , which encompasses 1515 genes and 2719 metabolic reactions of the Escherichia coli (E. coli) strain K-12 MG1655, a versatile organism for metabolic engineering to produce specific compounds.
@@
Figure 1: $BMLP_{active}$ encodes the GEM iML1515 as Boolean matrices and predicts the auxotrophic mutant phenotypes. It actively consults a data source to request ground truth labels, which minimises the expected experimental cost based on a user-provided cost function. The underlying BMLP iteratively refutes gene function annotation hypotheses inconsistent with labelled training examples. While $BMLP_{active}$ to date only learns from synthetic data, a laboratory robot can be integrated to perform selected experiments.
@@
Despite the extensive mapping of the E. coli genome to metabolic functions in iML1515, inaccurate phenotype predictions have been identified due to erroneous gene function annotations bernstein_evaluating_2023 . Learning gene function annotations in GEMs is computationally and empirically challenging. We built the first logic programming system on GEMs $BMLP_{active}$ (Fig. 1) based on our novel Boolean Matrix Logic Programming (BMLP) approach. $BMLP_{active}$ uses Boolean matrices to encode the biochemical and genetic relationships in iML1515 and enables high-throughput logical inferences to predict auxotrophic mutant phenotypes. $BMLP_{active}$ implements active learning to select auxotrophic mutant experiments that minimise the expected experimental cost via a user-defined cost function. $BMLP_{active}$ successfully learned gene function annotations through active learning while reducing 90% of the optional nutrient substance cost required by randomly selected experiments. It additionally reduced the number of experimental data needed compared to random experiment selection. When operating within a finite budget, our approach could deliver optimal experimental outcomes, whereas studies employing random experiment selection might not reach completion.
@@
Furthermore, we applied $BMLP_{active}$ to learning digenic function annotations, which holds significant implications for drug development and therapeutic interventions costanzo_global_2016 . Digenic functions are related to isoenzymes - two enzymes responsible for the same reaction. Digenic interactions between complex genotypes and phenotypes remains largely unexplored in most organisms costanzo_global_2019 , and their dynamics depend on the growth conditions bernstein_evaluating_2023 . A comprehensive understanding of digenic interactions would expedite strain engineering efforts. We show that $BMLP_{active}$ converged to the correct gene-isoenzyme mapping with as few as 20 training examples. This represents a significant promise of $BMLP_{active}$ to address complex genetic interactions on the whole genome level.
2 Representing a GEM using datalog
We refer the readers to LP_book ; ILP_foundation for terminology on datalog programs. A recursive program has a body literal that appears in the head of a clause. Specifically, we focused on linear and immediately recursive datalog as simple recursive datalog programs. A linear and immediately recursive datalog program has a single recursive clause where the recursive predicate appears in the body only once ioannidis_computation_1986 . This subset of recursive datalog programs can represent the relational structure of reaction pathways:
$pathway(X,Y)\leftarrow reaction(X,Y).$
$pathway(X,Y)\leftarrow reaction(X,Z),pathway(Z,Y).$
Petri nets are a class of directed bipartite graphs reisig_understanding_2013 , and they are a commonly used graphical tool to model processes in biological systems sahu_advances_2021 . we refer readers to reisig_understanding_2013 ; rozenberg_elementary_1998 for background on Petri net. Petri nets contain nodes marked by tokens (black dots), which indicate the availability of resources such as chemical metabolites. The presence of reactants or products determines which reaction pathways are viable and how the resources are allocated. We examined a class of Petri nets called the one-bounded elementary net (OEN) reisig_understanding_2013 , which allocates at most one token per node (definitions in Appendix A). We use OEN as a conceptual model of essential metabolites and pathways in a metabolic network. We provided an example of a simple metabolic network encoded as a linear and immediately recursive datalog program $\mathcal{P}_{1}$.
$\mathcal{P}_{1}:\left\{\begin{aligned} reaction\_1(c1\_c2,c3\_c4).&\\ reaction\_1(c1\_c2,c3).&\\ reaction\_1(c1\_c2,c4).&\\ reaction\_2(c3\_c4,c5).&\\ reaction\_3(c5,c4).&\\ reaction(X,Y)\leftarrow reaction\_1(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_2(X,Y).&\\ reaction(X,Y)\leftarrow reaction\_3(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Y).&\\ pathway(X,Y)\leftarrow reaction(X,Z),&\\ pathway(Z,Y).&\end{aligned}\right\}$
Each ground fact in the program describes a reaction between a set of reactants and a set of products. Evaluating the recursive program $pathway(X,Y)$ is equivalent to finding the union of token distributions that respect the viable reactions (see Appendix B).
3 Boolean matrix logic programming
We propose the Boolean Matrix Logic Programming (BMLP) approach, which uses Boolean matrices to evaluate datalog programs in contrast to the traditional symbolic logic program evaluation.
Definition 1 (Boolean Matrix Logic Programming (BMLP) problem)
A linear and immediately recursive datalog program $\mathcal{P}$ can be written as matrices for bottom-up evaluation ceri_datalog_1989 and evaluated by linear equations sato_linear_2017 . We created a BMLP algorithm [^1] At the time of submission, the algorithm had been implemented in SWI-Prolog (version 9.2). called iterative extension (BMLP-IE) (Fig. 2) to evaluate the datalog encoding of the metabolic network. BMLP-IE uses a combination of binary operations: Boolean matrix addition (ADD), multiplication (MUL) and equality (EQ) copilowish_matrix_1948 . This algorithm uses $O(n^{2})$ binary operations given $n$ total metabolites in the metabolic network (Appendix D).
Figure 2: Iterative extension. (a) The vector v encodes source chemical metabolites. All reactions are represented in the boolean matrices $\textbf{R}_{1}$ and $\textbf{R}_{2}$. (b) BMLP-IE computes $\textbf{v}^{*}$, the complete set of producible metabolites.
4 Active learning
4.1 Hypothesis compression
We use the compression score of a hypothesis $h$ to optimise the posterior probability. A learner can derive the relative frequency $p(h)$ for $h$ to be chosen based on the encoding of $h$. The compactness of $h$’s encoding reflects its prior probability $p(h)=2^{-size(h)}$ shannon_mathematical_1963 . We can use the compression function to compute the prior $p(h)$. Given a set of training examples $E$, the compression of a hypothesis progol ; ase_progol can be interpreted as
$compression(h,E)=size(E)-MDL_{h,E}$
$\frac{p(h|E)}{p(E|E)}=2^{compression(h,E)}$
We can rearrange this to obtain the normalised posterior probability ase_progol :
The posterior probability is maximal when compression is maximal. By searching for a hypothesis with maximum posterior probability, a learner can maximise its expected predictive accuracy progol . To learn an accurate hypothesis, the goal is to find a hypothesis with the highest compression score.
The compression function in $BMLP_{active}$ follows the Minimal Description Length (MDL) principle conklin_complexity-based_1994 . Since we have a fixed set of hypotheses without an explicit prior distribution, MDL allows us to make minimum assumptions about their prior. The most probable hypothesis $h$ for the training examples $E$ should be compact and have the fewest disagreements between $E$ and predictions made by $h$. This minimises $L(h)+L(E|h)$, where $L(h)$ is the descriptive complexity of $h$ and $L(E|h)$ is the descriptive complexity of encoding $E$ using $h$. We consider a compression function from ase_progol based on the MDL principle:
$compression(h,E)=|E^{+}|-\frac{|E^{+}|}{pc_{h}}(size(h)+fp_{h})$
where $E^{+}$ is the set of positive examples seen by the active learner, $pc_{h}$ is the number of positive predictions by $h$, and $fp_{h}$ is false positives covered by $h$. In other words, this compression function favours a general hypothesis with high coverage and penalises over-general hypotheses that incorrectly predict negative examples. The generality of $h$ can be estimated by computing its coverage $pc_{h}$ for a set of unlabelled instances progol . We look at all available experiment instances to obtain a good estimation of the generality of hypotheses. Our BMLP approach can facilitate efficient inference for this estimation.
4.2 Expected cost of experiments
$EC(\emptyset,T)=0$
$EC({h},T)=0$
$EC(H,T)=min_{t\in T}\,[C_{t}+p(t)EC(H_{t},T-\{t\})+(1-p(t))EC(\overline{H_{t}},T-\{t\})]$
$+(1-p(t))(mean_{t^{\prime}\in T-\{t\}}C_{t^{\prime}})J_{\overline{H_{t}}}]$
where $H_{t}$ and $\overline{H_{t}}$ are subsets of hypotheses $H$ consistent and inconsistent with $t$’s label. $p(t)$ is the probability that the outcome of the experiment $t$ is positive and $J_{H}=-\sum_{h\in H}p^{\prime}(h|E)\,log_{2}(p^{\prime}(h|E))$. The probability $p^{\prime}(h|E)$ is calculated from the compression function. To estimate $p(t)$, we compute the sum of the probabilities of the hypotheses consistent with a positive outcome of $t$. Users can flexibly define the experiment cost function $C_{t}$.
4.3 Active learning sample complexity
For some hypothesis space $H$ and background knowledge $BK$, let $V_{s}$ denote the version space of hypotheses consistent with $s$ training examples. For an active version space learner that selects one instance per iteration, $|V_{s}|$ denotes the size of the version space at the iteration $s$ of active learning. The shrinkage of the hypothesis space can be represented by the reduction ratio $\frac{|V_{s+1}|}{|V_{s}|}$ after querying the $s+1$ label hocquette_how_2018 . The minimal reduction ratio $p(x_{s+1},V_{s})$ is the minority ratio of the version space $V_{s}$ partitioned by an instance $x_{s+1}$.
Definition 2
(Minimal reduction ratio hocquette_how_2018 ) The minimal reduction ratio over the version space $V_{s}$ by sampled instance $x_{s+1}$ is
$p(x_{s+1},V_{s})=\frac{min(\,|\{h\in V_{s}|\,h\cup BK\models x_{s+1}\}|,|\{h\in V_{s}|\,h\cup BK\not\models x_{s+1}\}|)}{|V_{s}|}$
The minimal reduction ratio of an actively selected instance can be computed before it is labelled. While in reality training examples might not be as discriminative, the optional selection strategy is to select instances with minimal reduction ratios as close as possible to $\frac{1}{2}$ with the ability to eliminate up to 50% of the hypothesis space. We describe the sample complexity advantage of active learning over random example selection by the following bound on an active version space learner’s sample complexity. A passive learner is a learner using random example selection and it does not have control over the training examples it uses. Theorem 4.1 (proof in Appendix E) says that the number of instances needed to select by active learning should be some factor smaller than the number of randomly sampled examples given some predictive accuracy level.
Theorem 4.1 (Active learning sample complexity bound)
$s_{active}\leq\frac{\epsilon}{\epsilon+\phi}s_{passive}+c$ (1)
Given a desirable predictive error $\epsilon$, the factor in Theorem 4.1 can be estimated from the minimal reduction ratio $\phi$ over the instance space. When selected instances have a larger minimal reduction ratio, the sample complexity gain would be more significant. On the other hand, if instances have low discriminative power and the minimal reduction ratio tends to zero, the sample complexity gain would be minimal. In Section 6, we empirically show evidence supporting this theorem for $BMLP_{active}$ and randomly sampled experiments when learning gene function annotations in a GEM.
5 Implementation
Algorithm 1 $BMLP_{active}$
Output: A hypothesis $h_{max}\in H$.
1:  Let $T_{selected}=\emptyset$, $V_{0}=H$
2:  while $cost(T_{selected})\leq C$ and $|V_{j}|>1$ and $|T_{selected}|<N$ do
3:     if $T_{selected}==\emptyset$ then
4:        Select $t_{j}$ with the largest minimum reduction ratio $p(t_{j},V_{j})$
5:           from remaining experiments in $T$ to add to $T_{selected}$
6:     else
8:     end if
9:     break if $cost(T_{selected})>C$
10:     Consult the label $l$ of $t_{j}$ from the oracle $O$
12:     $V_{j+1}=\{h_{i}\,|\,h_{i}\in V_{j}\text{ and }\textbf{R}_{i,j}=l\}$
13:     $h_{max}=\mathrm{argmax}_{h_{i}\in V_{j+1}}\,compression(h_{i},T_{selected})$
14:  end while
The input background knowledge is created from a GEM model, which is typically accessible from GEM repositories king_bigg_2016 ; li_gotenzymes_2023 and libraries schellenberger_quantitative_2011 ; ebrahim_cobrapy_2013 . The user provides a set of candidate experiments, hypothesis candidates and an experiment cost function. Labels of experiment instances are requested from a data source, e.g. a laboratory or an online dataset. Labelled experimental data are considered ground truths. Initially, we randomly shuffle all candidate experiments and select an experiment with the most discriminative power if its cost is under the budget. Alternatively, a discriminative experiment with the lowest cost can be selected for a fixed initialisation.
Active learning in $BMLP_{active}$ is described by Algorithm 1. A classical method to address errors in metabolic pathways involves auxotrophic growth experiments, where specific genes are deleted to render the organism incapable of synthesising essential compounds beadle_genetic_1941 . We obtain binary phenotypic classifications of cell growth comparable to an unedited wild-type strain. We identified all essential metabolites in the wild-type strain from the BMLP-IE output. A positive label (1) is a phenotypic effect, and a negative label (0) is no phenotypic effect. We predict labels from the combinations of candidate experiments $T$ and hypotheses $H$ and store them in a Boolean matrix R of size $|H|\times|T|$. Given the j-th actively selected instance $t_{j}\in T$ and $h_{i}\in V_{j}\subseteq H$, $(\textbf{R})_{i,j}=1$ if $h_{i}\cup BK\models t_{j}$ and otherwise $(\textbf{R})_{i,j}=0$.
In each active learning cycle in $BMLP_{active}$, hypotheses inconsistent with ground truth experimental outcomes are pruned. $BMLP_{active}$ selects experiments to minimise the expected value of a user-defined cost function. This avoids repetitive predictions in hypothesis pruning and compression calculation. When computing hypothesis compression, we estimate the generality of a hypothesis by calculating its coverage for all unlabelled instances based on R. Currently, we only consider labels from synthetic or online phenotype data. However, $BMLP_{active}$ can be coupled with a high-throughput experimental workflow to automate experiments.
6 Experiments
6.1 Experiment 1: BMLP-IE runtime and predictions
Single thread (seconds) 20 CPUs (seconds)
SWI-Prolog $37.842\pm 9.668$ $5.089\pm 0.119$
SWI-Prolog + BMLP-IE $\mathbf{0.220\pm 0.099}$ $\mathbf{0.061\pm 0.008}$
Table 1: Mean runtime of 100 predictions in CPU time. BMLP-IE leads to a 170 times improvement in prediction time efficiency. Multi-threading BMLP-IE enhances runtime efficiency 600 times compared to base SWI-Prolog.
Here we demonstrate [^2] All experiments are performed on a workstation with Intel(R) Core(TM) i9-7900X CPU 3.30GHz (20 CPUs). that a SWI-Prolog application can be made significantly faster by involving Boolean matrix computation. We randomly sampled batches of 100 experiments and computed the average wall time from 10 repeats (Appendix G.2). Table 1 shows a 170 times improvement in prediction time by BMLP-IE compared to just using just SWI-Prolog. BMLP-IE uses only $\frac{1}{600}$ of the prediction time compared to predictions with just SWI-Prolog. This result demonstrates that BMLP-IE significantly improves the speed in predicting phenotypes from a genome-scale model.
Figure 3: Normalised confusion matrices. The predictive accuracy of BMLP-IE is $88.7\%$ and standard FBA is $93.4\%$. Each number has been normalised by the size of the experiment dataset in iML1515 .
FBA palsson_systems_2015 is widely used in the literature to quantitatively predict the cell growth rate based on fluxes of biomass metabolites at the steady state. We evaluated BMLP-IE’s phenotypic predictions against FBA based on the GEM model iML1515 iML1515 . We created normalised confusion matrices (Fig. 3) for BMLP-IE and the FBA predictions with respect to experimental data iML1515 . The confusion matrices show a comparable predictive accuracy by BMLP-IE to that of standard FBA. Compared with FBA’s growth rate predictions, BMLP-IE mainly misclassified reduced cell growth. BMLP-IE is easily adaptable when GEM is extended, but FBA requires extensive manual parameter updates in order to tune metabolic fluxes.
Figure 4: Total experimentation cost needed for the recovered models to reach different levels of predictive accuracy. The cost definitions and details are in Appendix G and H). The predictive accuracy of an empty hypothesis is 17.4%.
[…]
Fig. 4 shows that $BMLP_{active}$ spends 10% of the cost used by random experiment sampling when converging to the correct hypothesis. The y-axis is the total experimental reagent cost in $log_{10}$ calculated from selected experiments and normalised by the cost of the cheapest optional nutrient. A predictive accuracy of 100% indicates successful recovery of deleted gene function annotations. The correct annotations were recovered by $BMLP_{active}$ and random experiment selection with $10^{1.99}$ and $10^{2.99}$ experimental reagent costs. Given the current hypothesis space, experimental design space and our assumption that one gram per reagent was used, learning a single annotation required £$\pounds 3.8$ and £$\pounds 38$ experimental costs from $BMLP_{active}$ and random sampling. Consideration of differential resource factors in the user-defined experimental cost function could significantly increase the total experimental cost.
In addition, Fig. 5 shows that $BMLP_{active}$ reduces the number of labelled data to learn accurate gene function annotations, compared to random experiment selection. While random experiment selection requires 25 experiments to recover the deleted annotations, $BMLP_{active}$ only needed 3 experiments. Fig. 4 and Fig. 5 show $BMLP_{active}$ can simultaneously optimise a user-defined experimental cost function and reduce the total number of experiments to learn gene function annotations accurately. $BMLP_{active}$ is highly flexible and could be tailored for specific experimental objectives in a discovery process.
Figure 5: Reduction of the number of experiments needed to recover gene function annotations (sample complexity reduction). $BMLP_{active}$ learns fully accurate gene function annotations with 3 experiments, while random sampling requires 25 experiments to do so. The predictive accuracy of an empty hypothesis is 17.4 %. Predictive accuracy is higher for both methods initially since experiment selection is no longer constrained by reagent costs.
Fig. 6 shows the relationships between predictive error and the number of experiments in Theorem 4.1. $BMLP_{active}$ selected experiments with minimal reduction ratio $\phi\geq 0.5\%$. The ratio $\frac{\epsilon}{s_{active}}$ is greater than $\frac{\epsilon+\phi}{s_{passive}}$ for error $\epsilon>0$. There are fewer points on the active learning ratio curve since $BMLP_{active}$ converges to accuracy 100% using fewer training examples. The two curves intersect at $\epsilon=0$ since both converged to 100% accuracy. Theorem 4.1 shows that active learning ($BMLP_{active}$) can reduce the requirement for training examples compared with passive learning (random experiment selection) to achieve a target learning performance. In the context of scientific discovery, the experimental cost to arrive at a finding can be reduced since each actively selected experiment is more informative.
6.3 Experiment 3: learning digenic functions
We show the recovery of a known gene-isoenzyme association, which presents a bigger empirical challenge due to the need to explore a larger experimental design space. We focused on the aromatic amino acid biosynthesis pathway. It is a critical pathway for microorganism survival that contains important gene targets for engineering aromatic amino acid production strains. Although this pathway has been intensively studied for decades, gaps remain in the understanding of gene-gene relations price_filling_2018 . Fig. 7 shows the result of recovering tyrB’s isoenzyme function. We observed that $BMLP_{active}$ successfully recovers the correct tyrB isoenzyme function in 20 experiments. This demonstrates $BMLP_{active}$’s ability to discover gene-isoenzyme associations in GEM. However, to learn all isoenzyme functions, a high-throughput experimentation procedure is required and we further discuss this in Section 8.
$BMLP_{active}$ significantly reduced the number of experiments needed compared to random experiment selection. $BMLP_{active}$ provides higher information gain from each experiment and can guarantee recovery with as few as 20 experiments. While random sampling could prune most candidate hypotheses, it failed to eliminate competitive hypotheses further. In contrast to $BMLP_{active}$, random sampling could not recover this isoenzyme function with 64 experiments and could only do this occasionally with more than 10 times the number of experiments required by $BMLP_{active}$. Although we used a reduced set of genes in this demonstration, the combinatorial space was sufficiently large that random experiment selection became non-viable as an experimentation strategy in a discovery process. The remarkable increase in efficiency with $BMLP_{active}$ with this task demonstrates it has potential even as the experimental design space grows.